OpenStudy (anonymous):

Write as sine, cosine, or tangent of an angle sin(pi/3) cos(pi/7) - sin(pi/7) cos(pi/3). And if you can, please explain how you got this

ganeshie8 (ganeshie8):

use this formula :- \(\large \sin (A-B) = \sin (A) \cos (B) - \cos (A) \sin (B)\)

OpenStudy (anonymous):

ok so i did sin(a-b) as sin(pi/3)-(pi/7) which gave me 4pi/21. Is that the answer? Please excuse my stupidity

ganeshie8 (ganeshie8):

thats is \(\large \color{red}{\checkmark}\) = sin(pi/3-pi/7) = sin(4pi/21)

OpenStudy (anonymous):

oh my godness thank you so much!!

ganeshie8 (ganeshie8):

np... u wlc (: good work !